# Jet transport coefficients by elastic and radiative scatterings in the strongly interacting quark-gluon plasma

### v1

@@ Abstract
We extend the investigation on jet transport coefficients within the effective Dynamical QuasiParticle Model (DQPM) – constructed for the description of non-perturbative QCD phenomena of the strongly interacting quark-gluon plasma (sQGP) in line with the lattice QCD equation-of-state – by accounting for inelastic $2\to 3$ reactions with gluon radiation additionally to the elastic scattering of partons. The elastic and inelastic reactions are calculated explicitly within leading-order Feynman diagrams with effective propagators and vertices from the DQPM by accounting for all channels and their interferences. We present the results for the jet transport coefficients such as the transverse momentum transfer squared $\hat{q}$ per unit length as well as the energy loss $\Delta E=dE/dx$ per unit length in the sQGP and investigate their dependence on the temperature $T$ and momentum of the jet parton depending on the choice of the strong coupling constant $\alpha_{s}$ in thermal, jet parton and radiative vertices. For the latter we consider different scenarios used in the literature and find a very strong dependence of $\hat{q}$ and $\Delta E$ on the choice of $\alpha_{s}$. Moreover, we explore the relation of $\hat{q}/T^{3}$ to the ratio of specific shear viscosity to entropy density $\eta/s$ and show that the ratio $T^{3}/\hat{q}$ to $\eta/s$ has a strong $T$ dependence – especially when approaching to $T_{c}$ – on the choice of $\alpha_{s}$ in scattering vertices.
@@ I Introduction
Jets are considered as one of the penetrating probes that allow to obtain information about the strong interactions in the QGP medium created in heavy-ion collisions. As observed experimentally at RHIC Adams et al. (2003); Adler et al. (2003) and at LHC Aamodt et al. (2011); Aad et al. (2010) jets, produced in these A+A collisions, are modified compared to those produced in proton-proton (p+p) collisions due to the scattering of partons from the jet shower with the partons from the QGP environment. Theoretical understanding of the experimental data on jet attenuation in heavy-ion collisions stimulated a lot of a theoretical efforts on an understanding of jet properties in both equilibrium and non-equilibrium cases Thoma and Gyulassy (1991); Dokshitzer et al. (1991); Mrowczynski (1991); Gyulassy and Wang (1994); Baier et al. (1997a, b, 1998); Zakharov (1996); Gyulassy et al. (2000); Wiedemann (2000); Guo and Wang (2000); Arnold et al. (2002); Armesto et al. (2004); Moore and Teaney (2005); Djordjevic and Heinz (2008); Zapp et al. (2009a); Majumder (2012); Caron-Huot and Gale (2010); Stojku et al. (2021); Barata and Mehtar-Tani (2020); Cao and Wang (2021). The jet energy loss in the QGP medium occurs due to elastic $2\to 2$ partonic scatterings as well as by radiative $2\to 3$ processes with the emission of a gluon. The soft gluon radiation is screened in the medium due to the coherence effect called Landau-Pomeranchuk-Migdal (LPM) effect Landau and Pomeranchuk (1953); Migdal (1956) which has to be accounted for the interpretation of experimental data. The first calculations of the gluon Bremsstrahlung processes with the non-Abelian LPM effect in the context of perturbative QCD Baier et al. (1995, 1996); Wang et al. (1995); Zakharov (1996, 2021) showed a suppression of soft gluon radiation due to the destructive interference of the emitted gluons in the medium contrary to the vacuum, cf. also Refs. Gyulassy et al. (2000); Wiedemann (2000); Arnold (2009); Zapp et al. (2011); Armesto et al. (2012); Casalderrey-Solana and Iancu (2011); Ke et al. (2019); Senzel (2020). A further formulation of radiative energy loss - as a dynamical process in the QGP medium of the thermal scattering centers - has been derived by Arnold, Moore, and Yaffe (AMY) within thermal field theory Arnold et al. (2003, 2002, 2000) and developed further in recent years by going beyond the limitations of soft-emission approximations Arnold et al. (2020, 2022). Also progress has been achieved in the resummation of multiple scatterings Andres et al. (2022, 2023); Barata and Mehtar-Tani (2020).
@@
Since the jets are multi-scale objects, the microscopic simulation of jet propagation and interactions with the QCD medium is notoriously difficult. The description of partonic energy loss in the QGP is widely studied in the literature in terms of transport coefficients such as the transverse momentum transfer squared $\hat{q}$ per unit length to characterize the jet transverse momentum broadening in the medium as well as the energy loss $\Delta E=dE/dx$ per unit length to account for the modification of jet energy during propagation in the medium.
@@
The $\hat{q}$ represents a measure of the interaction between a high-energetic jet and the QGP medium, which is defined as the amount of momentum transfer that the hard parton experiences per unit length as it travels through the dense QGP medium. The value of this coefficient depends on several factors, including the coupling strength of the interaction, the nature of the plasma (whether it is dominated by quasi-particles or not), and the underlying micro-physics of many-body QCD matter.
@@
Recently, in Ref. Grishmanovskii et al. (2022) the jet transport coefficients $\hat{q}$ and $\Delta E$ have been studied for elastic $2\to 2$ interactions of partons within the effective dynamical quasi-particle model (DQPM) constructed to describe the non-perturbative properties of the sQGP at finite temperature $T$ and baryon chemical potential $\mu_{B}$ in terms of strongly interacting off-shell partons (quarks and gluons) with dynamically generated spectral functions whose properties are adjusted to reproduce the lQCD EoS for the QGP in thermodynamic equilibrium. It has been shown that $\hat{q}$ and $\Delta E$ show a strong temperature dependence due to the growing strong coupling $\alpha_{s}(T)$ when approaching the critical temperature $T_{c}$. However, additionally, to elastic scattering, the gluon radiative processes $2\to 3$ are also important for an understanding of the high energetic jet attenuation in heavy-ion collisions, whose contribution grows with increasing energy of jet partons - cf. Djordjevic (2006); Stojku et al. (2021); Cao and Wang (2021) and Refs. therein.
@@
In the present work we aim to understand how important gluon radiative reactions are for the jet transport coefficient in the case of massive quarks and gluons and explore the temperature and momentum dependence of the transport coefficients. The latter can be of great interest in the case of jet quenching models such as JEWEL Zapp et al. (2009b), JET Burke et al. (2014), JETSCAPE Cao et al. (2021) etc. It is important to note that we do not use the assumption of perturbative QCD (pQCD) that the interaction between an energetic parton and the medium is dominated by small-angle scattering and induced gluon radiation. In this work, we aim to evaluate jet transport coefficients taking into account properly all diagrams and channels involved in $2\to 2$ and $2\to 3$ partonic scattering with massive medium partons as described by DQPM, i.e. the in-medium modification of jet partons occurs by sequences of independent elastic and inelastic reactions with thermal partons from the medium at given temperature $T$.
The framework of quasiparticle models facilitates its implementation in the transport model for the dynamical evolution of the QGP matter. In particular, the off-shell quasi-particle model DQPM is implemented in the Parton-Hadron-String Dynamics (PHSD) transport approach Cassing and Bratkovskaya (2009); Moreau et al. (2019), whereas the on-shell QPM Plumari et al. (2011) is implemented in the Catania transport approach Scardina et al. (2017).
In this study we also investigate the dependence of elastic and inelastic transport coefficients on the choice of the strong coupling constant $\alpha_{s}$ used in thermal, jet parton and gluon radiative vertices, which could differ from the coupling constant for the thermalized partons since the jet parton is not in equilibrium with the QGP medium Zakharov (2023); Boguslavski et al. (2023); Shi et al. (2023). For that goal we select 4 models for $\alpha_{s}$ – used in the literature – and compare the results with default the DQPM $\alpha_{s}(T)$. Also we explore the dependence of elastic and inelastic $\hat{q}$ and $dE/dx$ on the mass of the emitted gluon. Some model cases we compare to the results from the BAMPS (Boltzmann Approach to Multi-Parton Scatterings) pQCD cascade which explicitly includes $2\to 2$ and $2\to 3$ (as well as backward) reactions Uphoff et al. (2015); Fochler et al. (2013) and the LPM effect Senzel (2020).
Furthermore, we investigate the relation of $\hat{q}/T^{3}$ to the ratio of specific shear viscosity to entropy density $\eta/s$ discussed in Müller (2021).
This paper is organized as follows. In Section II we recall the main ideas of the DQPM, in Section III we describe the framework for the calculation of transport coefficients and present the scenarios for the coupling constant. In Sections IV, V and VI we report on the results for the radiative differential cross sections and transport coefficients. We summarize our study in Section VII.
@@ II Dynamical Quasiparticle Model
The Dynamical Quasiparticle Model (DQPM) Peshier and Cassing (2005); Cassing (2007a, b); Berrehrah et al. (2016a); Moreau et al. (2019); Soloveva et al. (2021) is an effective model which describes the Quark-Gluon Plasma (QGP) in terms of strongly interacting quarks and gluons. This approach involves fitting the properties of these quasiparticles to match the results of lattice QCD calculations in thermal equilibrium and at vanishing chemical potential. Here we briefly recall the basic ideas of the DQPM. The quasi-particles in the DQPM are characterized by ”dressed” propagators, i.e. single-particle (two-point) Green’s functions, which have the form
$G^{R}_{j}(\omega,{\bf p})=\frac{1}{\omega^{2}-{\bf p}^{2}-M_{j}^{2}+2i\gamma_{j}\omega}$ (1)
for quarks, antiquarks, and gluons ($j=q,\bar{q},g$), using $\omega=p_{0}$ for energy, the widths $\gamma_{j}$ and the masses $M_{j}$ and the complex self-energies for gluons $\Pi=M_{g}^{2}-2i\omega\gamma_{g}$ and for (anti)quarks $\Sigma_{q}=M_{q}^{2}-2i\omega\gamma_{q}$, where the real part of the self-energies is associated with dynamically generated thermal masses, while the imaginary part provides information about the lifetime and reaction rates of the particles.
@@
$\rho_{j}(\omega,{\bf p})$ $=\frac{\gamma_{j}}{\tilde{E}_{j}}\left(\frac{1}{(\omega-\tilde{E}_{j})^{2}+\gamma^{2}_{j}}-\frac{1}{(\omega+\tilde{E}_{j})^{2}+\gamma^{2}_{j}}\right)$
$\equiv\frac{4\omega\gamma_{j}}{\left(\omega^{2}-\mathbf{p}^{2}-M^{2}_{j}\right)^{2}+4\gamma^{2}_{j}\omega^{2}}$ (2)
with $\tilde{E}_{j}^{2}({\bf p})={\bf p}^{2}+M_{j}^{2}-\gamma_{j}^{2}$. The spectral function is antisymmetric in $\omega$ and normalized as
@@
The DQPM involves a model ansatz for the masses $M_{j}(T,\mu_{q})$ and widths $\gamma_{j}(T,\mu_{q})$ as functions of the temperature $T$ and the quark chemical potential $\mu_{q}$. By comparison of the entropy density – computed within the DQPM framework – to the lQCD data, one can fix the few parameters used in the ansatz for the quasiparticle masses and widths. The following ansatz is used in the DQPM for the definition of the quasiparticle properties (masses and widths) as functions of $T$ and $\mu_{q}$. The dynamical quasiparticle pole masses are given by the HTL thermal mass in the asymptotic high-temperature regime - cf. Bellac (2011); Linnyk et al. (2016)
@@
where $N_{c}\ (=3)$ stands for the number of colors and $N_{f}\ (=3)$ denotes the number of light flavors. Eq. (5) determines pole masses for the ($u,d$) quarks, the strange quark has a larger bare mass for controlling the strangeness ratio in the QGP. Empirically, we find $M_{s}(T,\mu_{B})=M_{u/d}(T,\mu_{B})+\Delta M$, where $\Delta M\simeq 30$ MeV has been fixed once in comparison to experimental data Moreau et al. (2019).
@@
Here $c_{m}=14.4$ is related to a magnetic cutoff, which is an additional parameter in the DQPM, while $C_{q}=\dfrac{N_{c}^{2}-1}{2N_{c}}=4/3$ and $C_{g}=N_{c}=3$ are the QCD color factors for quarks and gluons, respectively. We also assume that all (anti-)quarks have the same $T$-dependence for the width.
The coupling constant $g^{2}=4\pi\alpha_{s}$ essentially defines the masses and widths of quasiparticles - cf. Eqs. (5), (6). The thermal properties of quasiparticles and their interactions (observed via transport coefficients) thus strongly depend on the coupling constant $g$.
In the DQPM $g^{2}$ is extracted from lattice Quantum Chromodynamics (lQCD) data on the entropy density $s$ by a parametrization method introduced in Ref. Berrehrah et al. (2016b). There it has been found that for a given value of $g^{2}$, the ratio $s(T,g^{2})/T^{3}$ is almost constant for different temperatures, i.e. ${\frac{\partial}{\partial T}}(s(T,g^{2})/T^{3})=0$. Therefore the entropy density $s$ and the dimensionless equation of state in the DQPM is a function of the effective coupling only, i.e. $s(T,g^{2})/s_{SB}(T)=f(g^{2})$ where $s^{QCD}_{SB}=19/9\pi^{2}T^{3}$ is the Stefan-Boltzmann entropy density. Thus, by inverting the $f(g^{2})$ function, the coupling constant $g^{2}$ can be directly obtained from the parametrization of lQCD data for the entropy density $s(T,\mu_{B}=0)$ at zero baryon chemical potential:
$g^{2}(T,\mu_{B}=0)=d\left((s(T,0)/s^{QCD}_{SB})^{e}-1\right)^{f}.$ (7)
Here $d=169.934,e=-0.178434$ and $f=1.14631$ are the dimensionless parameters obtained by adjusting the quasiparticle entropy density $s(T,\mu_{B}=0)$ to the lQCD data provided by the BMW Collaboration Borsanyi et al. (2012, 2014).
The DQPM $\alpha_{s}$ accounts for nonperturbative effects which is larger compared to the analytical 2- or 1-loop running constant Caswell (1974) approaching low temperatures.
The extension of the coupling constant to finite baryon chemical potential $\mu_{B}$ is realized using a scaling hypothesis Cassing (2009) that works up to $\mu_{B}\approx 500$ MeV.
Thus, the DQPM provides the quasiparticle properties, dressed propagators, and coupling constant, which can be used to evaluate the scattering amplitudes as well as the cross sections and the transport coefficients of quarks and gluons in the QGP as a function of the temperature and the chemical potential - cf. Refs. Berrehrah et al. (2014a); Moreau et al. (2019); Grishmanovskii et al. (2023).

### v2

@@ Abstract
We extend the investigation of jet transport coefficients within the effective Dynamical QuasiParticle Model (DQPM) – constructed to describe nonperturbative QCD phenomena of the strongly interacting quark-gluon plasma (sQGP) in line with the lattice QCD equation of state – by accounting for inelastic $2\to 3$ reactions with gluon radiation in addition to the elastic scattering of partons. The elastic and inelastic reactions are calculated explicitly within leading-order Feynman diagrams with effective propagators and vertices from the DQPM by accounting for all channels and their interferences. We present the results for the jet transport coefficients such as the transverse momentum transfer squared $\hat{q}$ per unit length as well as the energy loss $dE/dx$ per unit length in the sQGP and investigate their dependence on the temperature $T$ and momentum of the jet parton depending on the choice of the strong coupling constant $\alpha_{s}$ in thermal, jet parton and radiative vertices. For the latter, we consider different scenarios used in the literature and find a very strong dependence of $\hat{q}$ and $dE/dx$ on the choice of $\alpha_{s}$. Moreover, we explore the relation of $\hat{q}/T^{3}$ to the ratio of specific shear viscosity to entropy density $\eta/s$ and show that the ratio $T^{3}/\hat{q}$ to $\eta/s$ has a strong $T$ dependence – especially when approaching $T_{c}$ – on the choice of $\alpha_{s}$ in scattering vertices.
@@ I Introduction
Jets are considered as one of the penetrating probes that allow to obtain information about the strong interactions in the QGP medium created in heavy-ion collisions. As observed experimentally at RHIC Adams et al. (2003); Adler et al. (2003) and at LHC Aamodt et al. (2011); Aad et al. (2010), jets produced in these A+A collisions are modified compared to those produced in proton-proton (p+p) collisions due to the scattering of partons from the jet shower with the partons from the QGP environment. Theoretical understanding of the experimental data on jet attenuation in heavy-ion collisions stimulated a lot of theoretical efforts on an understanding of jet properties in both equilibrium and non-equilibrium cases Thoma and Gyulassy (1991); Dokshitzer et al. (1991); Mrowczynski (1991); Gyulassy and Wang (1994); Baier et al. (1997a, b, 1998); Zakharov (1996); Gyulassy et al. (2000); Wiedemann (2000); Guo and Wang (2000); Arnold et al. (2002); Armesto et al. (2004); Moore and Teaney (2005); Djordjevic and Heinz (2008); Zapp et al. (2009a); Majumder (2012); Caron-Huot and Gale (2010); Stojku et al. (2021); Barata and Mehtar-Tani (2020); Cao and Wang (2021). The jet energy loss in the QGP medium occurs due to elastic $2\to 2$ partonic scatterings as well as by radiative $2\to 3$ processes with the emission of a gluon. The soft gluon radiation is screened in the medium due to the coherence effect called Landau-Pomeranchuk-Migdal (LPM) effect Landau and Pomeranchuk (1953); Migdal (1956), which has to be accounted for in the interpretation of experimental data. The first calculations of the gluon Bremsstrahlung processes with the non-Abelian LPM effect in the context of perturbative QCD Baier et al. (1995, 1996); Wang et al. (1995); Zakharov (1996, 2021) showed a suppression of soft gluon radiation due to the destructive interference of the emitted gluons in the medium, contrary to the vacuum; cf. also Refs. Gyulassy et al. (2000); Wiedemann (2000); Arnold (2009); Zapp et al. (2011); Armesto et al. (2012); Casalderrey-Solana and Iancu (2011); Ke et al. (2019); Senzel (2020). A further formulation of radiative energy loss – as a dynamical process in the QGP medium of the thermal scattering centers – has been derived by Arnold, Moore, and Yaffe (AMY) within thermal field theory Arnold et al. (2003, 2002, 2000) and developed further in recent years by going beyond the limitations of soft-emission approximations Arnold et al. (2020, 2022). Also, progress has been achieved in the resummation of multiple scatterings Andres et al. (2022, 2023); Barata and Mehtar-Tani (2020).
@@
Since the jets are multi-scale objects, the microscopic simulation of jet propagation and interactions with the QCD medium is notoriously difficult. The description of partonic energy loss in the QGP is widely studied in the literature in terms of transport coefficients, such as the transverse momentum transfer squared $\hat{q}$ per unit length to characterize the jet transverse momentum broadening in the medium, as well as the energy loss $dE/dx$ per unit length to account for the modification of jet energy during propagation in the medium.
@@
The $\hat{q}$ represents a measure of the interaction between a high-energetic jet and the QGP medium, which is defined as the amount of momentum transfer that the hard parton experiences per unit length as it travels through the dense QGP medium. The value of this coefficient depends on several factors, including the coupling strength of the interaction, the nature of the plasma (whether it is dominated by quasiparticles or not), and the underlying micro-physics of many-body QCD matter.
@@
Recently, in Ref. Grishmanovskii et al. (2022) the jet transport coefficients $\hat{q}$ and $dE/dx$ have been studied for elastic $2\to 2$ interactions of partons within the effective dynamical quasiparticle model (DQPM), constructed to describe the nonperturbative properties of the sQGP at finite temperature $T$ and baryon chemical potential $\mu_{B}$ in terms of strongly interacting off-shell partons (quarks and gluons) with dynamically generated spectral functions, whose properties are adjusted to reproduce the lQCD EoS for the QGP in thermodynamic equilibrium. It has been shown that $\hat{q}$ and $dE/dx$ show a strong temperature dependence due to the growing strong coupling $\alpha_{s}(T)$ when approaching the critical temperature $T_{c}$. However, additionally to elastic scattering, the gluon radiative processes $2\to 3$ are also important for an understanding of the high energetic jet attenuation in heavy-ion collisions, whose contribution grows with increasing energy of jet partons – cf. Refs. Djordjevic (2006); Stojku et al. (2021); Cao and Wang (2021) and references therein.
@@
In the present work, we aim to understand how important gluon radiative reactions are for the jet transport coefficient in the case of massive quarks and gluons and explore the temperature and momentum dependence of the transport coefficients. The latter can be of great interest in the case of jet quenching models, such as JEWEL Zapp et al. (2009b), JET Burke et al. (2014), JETSCAPE Cao et al. (2021), and etc. It is important to note that we do not use the assumption of perturbative QCD (pQCD), which states that the interaction between an energetic parton and the medium is dominated by small-angle scattering and induced gluon radiation. In this work, we aim to evaluate jet transport coefficients by properly taking into account all diagrams and channels involved in $2\to 2$ and $2\to 3$ partonic scattering with massive medium partons as described by DQPM, i.e., the in-medium modification of jet partons occurs by sequences of independent elastic and inelastic reactions with thermal partons from the medium at given temperature $T$.
The framework of quasiparticle models facilitates its implementation in the transport model for the dynamical evolution of the QGP matter. In particular, the off-shell quasiparticle model DQPM is implemented in the Parton-Hadron-String Dynamics (PHSD) transport approach Cassing and Bratkovskaya (2009); Moreau et al. (2019), whereas the on-shell QPM Plumari et al. (2011) is implemented in the Catania transport approach Scardina et al. (2017).
In this study, we also investigate the dependence of elastic and inelastic transport coefficients on the choice of the strong coupling constant $\alpha_{s}$ used in thermal, jet parton and gluon radiative vertices, which could differ from the coupling constant for the thermalized partons, since the jet parton is not in equilibrium with the QGP medium Zakharov (2023); Boguslavski et al. (2023); Shi et al. (2023). For that goal we select four models for $\alpha_{s}$ – used in the literature – and compare the results with the default DQPM $\alpha_{s}(T)$. Also we explore the dependence of elastic and inelastic $\hat{q}$ and $dE/dx$ on the mass of the emitted gluon. Some model cases we compare to the results from the Boltzmann Approach to Multi-Parton Scatterings (BAMPS) pQCD cascade, which explicitly includes $2\to 2$ and $2\to 3$ (as well as backward) reactions Uphoff et al. (2015); Fochler et al. (2013) as well as the LPM effect Senzel (2020).
Furthermore, we investigate the relation of $\hat{q}/T^{3}$ and the specific shear viscosity $\eta/s$ discussed in Müller (2021).
This paper is organized as follows. In Sec. II we recall the main ideas of the DQPM, in Sec. III we describe the framework for the calculation of transport coefficients and present the scenarios for the coupling constant. In Secs. IV, V, and VI we report on the results for the radiative differential cross sections and transport coefficients. We summarize our study in Sec. VII.
@@ II Dynamical Quasiparticle Model
The Dynamical Quasiparticle Model (DQPM) Peshier and Cassing (2005); Cassing (2007a, b); Berrehrah et al. (2016a); Moreau et al. (2019); Soloveva et al. (2021) is an effective model that describes the Quark-Gluon Plasma (QGP) in terms of strongly interacting quarks and gluons. This approach involves fitting the properties of these quasiparticles to match the results of lattice QCD calculations in thermal equilibrium and at vanishing chemical potential. Here we briefly recall the basic ideas of the DQPM. The quasiparticles in the DQPM are characterized by ”dressed” propagators, i.e., single-particle (two-point) Green’s functions, which have the form
$G^{R}_{j}(\omega,\mathbf{p})=\frac{1}{\omega^{2}-\mathbf{p}^{2}-M_{j}^{2}+2i\gamma_{j}\omega}$ (1)
for quarks, antiquarks, and gluons ($j=q,\bar{q},g$), using $\omega=p_{0}$ for energy, the widths $\gamma_{j}$ and the masses $M_{j}$, and the complex self-energies for gluons $\Pi=M_{g}^{2}-2i\omega\gamma_{g}$ and for (anti)quarks $\Sigma_{q}=M_{q}^{2}-2i\omega\gamma_{q}$, where the real part of the self-energies is associated with dynamically generated thermal masses, while the imaginary part provides information about the lifetime and reaction rates of the particles.
@@
$\rho_{j}(\omega,\mathbf{p})$ $=\frac{\gamma_{j}}{\tilde{E}_{j}}\left(\frac{1}{(\omega-\tilde{E}_{j})^{2}+\gamma^{2}_{j}}-\frac{1}{(\omega+\tilde{E}_{j})^{2}+\gamma^{2}_{j}}\right)$
$\equiv\frac{4\omega\gamma_{j}}{\left(\omega^{2}-\mathbf{p}^{2}-M^{2}_{j}\right)^{2}+4\gamma^{2}_{j}\omega^{2}},$ (2)
with $\tilde{E}_{j}^{2}(\mathbf{p})=\mathbf{p}^{2}+M_{j}^{2}-\gamma_{j}^{2}$. The spectral function is antisymmetric in $\omega$ and normalized as
@@
The DQPM involves a model ansatz for the masses $M_{j}(T,\mu_{q})$ and widths $\gamma_{j}(T,\mu_{q})$ as functions of the temperature $T$ and the quark chemical potential $\mu_{q}$. By comparison of the entropy density – computed within the DQPM framework – to the lQCD data, one can fix the few parameters used in the ansatz for the quasiparticle masses and widths. The following ansatz is used in the DQPM for the definition of the quasiparticle properties (masses and widths) as functions of $T$ and $\mu_{q}$. The dynamical quasiparticle pole masses are given by the HTL thermal mass in the asymptotic high-temperature regime – cf. Bellac (2011); Linnyk et al. (2016)
@@
where $N_{c}\ (=3)$ stands for the number of colors and $N_{f}\ (=3)$ denotes the number of light flavors. Equation (5) determines the pole masses for the ($u,d$) quarks; the strange quark has a larger bare mass for controlling the strangeness ratio in the QGP. Empirically, we find $M_{s}(T,\mu_{B})=M_{u/d}(T,\mu_{B})+\Delta M$, where $\Delta M\simeq 30$ MeV has been fixed once in comparison to experimental data Moreau et al. (2019).
@@
Here $c_{m}=14.4$ is related to a magnetic cutoff, which is an additional parameter in the DQPM, while $C_{q}=\dfrac{N_{c}^{2}-1}{2N_{c}}=4/3$ and $C_{g}=N_{c}=3$ are the QCD color factors for quarks and gluons, respectively. We also assume that all (anti)quarks have the same $T$ dependence for the width.
The coupling constant $g^{2}=4\pi\alpha_{s}$ essentially defines the masses and widths of quasiparticles – cf. Eqs. (5), (6). The thermal properties of quasiparticles and their interactions (observed via transport coefficients) thus strongly depend on the coupling constant $g$.
In the DQPM, $g^{2}$ is extracted from lattice Quantum Chromodynamics (lQCD) data on the entropy density $s$ by a parametrization method introduced in Ref. Berrehrah et al. (2016b). There it has been found that for a given value of $g^{2}$, the ratio $s(T,g^{2})/T^{3}$ is almost constant for different temperatures, i.e., ${\frac{\partial}{\partial T}}(s(T,g^{2})/T^{3})=0$. Therefore, the entropy density $s$ and the dimensionless equation of state in the DQPM is a function of the effective coupling only, i.e., $s(T,g^{2})/s_{SB}(T)=f(g^{2})$, where $s^{\mathrm{QCD}}_{SB}=19/9\pi^{2}T^{3}$ is the Stefan-Boltzmann entropy density. Thus, by inverting the $f(g^{2})$ function, the coupling constant $g^{2}$ can be directly obtained from the parametrization of lQCD data for the entropy density $s(T,\mu_{B}=0)$ at zero baryon chemical potential:
$g^{2}(T,\mu_{B}=0)=d\left((s(T,0)/s^{\mathrm{QCD}}_{SB})^{e}-1\right)^{f}.$ (7)
Here $d=169.934,e=-0.178434$, and $f=1.14631$ are the dimensionless parameters obtained by adjusting the quasiparticle entropy density $s(T,\mu_{B}=0)$ to the lQCD data provided by the BMW Collaboration Borsanyi et al. (2012, 2014).
The DQPM $\alpha_{s}$ accounts for nonperturbative effects and is larger compared to the analytical two- or one-loop running constant Caswell (1974) when approaching low temperatures.
The extension of the coupling constant to finite baryon chemical potential $\mu_{B}$ is realized using a scaling hypothesis Cassing (2009), which works up to $\mu_{B}\approx 500$ MeV.
Thus, the DQPM provides the quasiparticle properties, dressed propagators, and coupling constant, which can be used to evaluate the scattering amplitudes as well as the cross sections and the transport coefficients of quarks and gluons in the QGP as a function of the temperature and the chemical potential – cf. Refs. Berrehrah et al. (2014a); Moreau et al. (2019); Grishmanovskii et al. (2024).